OpenStudy (anonymous):

find the fractional equivalent of 0.41717....

OpenStudy (anonymous):

Try to write it as \[ \frac{4}{10}+\frac{17 \sum _{n=0}^{\infty } \frac{1}{10^{2 n}}}{10^3} \]

OpenStudy (anonymous):

\[0.41717=\frac{4}{10}+\frac{17}{10^3}+\frac{17}{10^5}+\frac{17}{10^7}+\cdots \]

OpenStudy (anonymous):

\[ 0.41717=\frac{4}{10}+\frac{17}{10^3}+\frac{17}{10^5}+\frac{17}{10^7}+\cdots=\\ \frac{4}{10}+\frac{17}{10^3}\left(1+\frac{1}{10^2}+\frac{1 }{10^4}+\cdots\right)= \]

OpenStudy (anonymous):

Use geometric series methods to show that \[ \sum _{n=0}^{\infty } \frac{1}{10^{2 n}}=\frac{100}{99} \]

OpenStudy (anonymous):

\[ 0.41717\cdots=\frac{4}{10}+\frac{17}{10^3}+\frac{17}{10^5}+\frac{17}{10^7}+\cdots=\\ \frac{4}{10}+\frac{17}{10^3}\left(1+\frac{1}{10^2}+\frac{1 }{10^4}+\cdots\right)= \frac{4}{10}+\frac{17}{10^3}\sum _{n=0}^{\infty } \frac{1}{10^{2 n}}=\\ \frac{4}{10}+\frac{17}{10^3}\frac{100}{99}=\frac {413}{990} \]

OpenStudy (raden):

maybe easier like this : let x=0.41717... (1) 10x=4.1717... (2) 1000x=417.17... (3) (3)-(2) : 990x=413 x=413/990